# AGRAG: Advanced Graph-based Retrieval-Augmented Generation for LLMs

### v1

@@ IV-A4 Hyperparameter and Hardware Settings
For our AGRAG, the length of text chunks $l_{t}$ is set to 256, and token overlap between chunks $l_{o}$ to 32, the number of raw mapped triples $k_{a}$ and the number of hybrid retrieved chunks $k_{r}$ are both set to 5, to align with the setting of HippoRAG2 and GraphRAG-bench. AGRAG’s unique hyperparameter, the entity extraction threshold, and the maximum n-gram, were set to 0.5 and 3, respectively. Please refer to subsection IV-E for further parameter sensitivity analysis.
All our experiments were carried out and evaluated with Qwen2.5-14B-Instruct [77] LLM backbone if not further illustrated. We apply the Contriever [78] embedding model for HippoRAG2 and AGRAG’s dense passage retrieval procedure, and bge-large-en-v1.5 [79] embedding model for other baselines’ LLM embedding. All experiments are conducted on an Intel(R) Xeon(R) Gold 5220R @ 2.20GHz CPU and a single NVIDIA A100-SXM4-80GB GPU.
@@ IV-B Effectiveness Evaluation
Other than Qwen-2.5-14B-Instruct, we also test the performance of AGRAG with GPT-4o [74] and the reinforcement learning based RAG reasoning backbone R1-Searcher-Qwen-2.5-7B [75] in Table IV. Due to cost and space considerations, we only compare AGRAG’s performance with the previous state-of-the-art model HippoRAG2 [31] on GraphRAG-bench benchmark [34] for GPT-4o, as the text classification task has a large number of texts to be classified, and can lead to high token cost (Over 7,000 USD on WOS dataset, \$15 for 1 million tokens). For R1-Searcher, we further compare NaiveRAG [15], as it is the RAG framework used in its original paper [75]. Based on the experiment results in Table III and Table IV, we have observations as follow:
Firstly, in line with recent findings [34, 35, 36], NaiveRAG [15] achieves better performance than all tested walking-based baselines on Contextual Summarization and Creative Generation tasks. This is because, walking-based baselines overlook the global, and coarse-grained information that is essential for effective summarization and creative generation. This result also suggests that the fine-grained, phrase-level information they capture is less beneficial for summarization compared to higher-level, coarse-grained insights, which help the LLM form a holistic understanding of the text chunks. Our AGRAG achieves the best performance on these two tasks by providing the reasoning subgrah, as well as the hybrid retrieved text chunks to the LLM, thereby offering more comprehensive coarse to fine-grained context for query answering.
Secondly, on simpler Complex Reasoning and Fact Retrieval tasks, the more recent walking-based models achieve the best performance. This indicates that the phrase-level dependencies, both long and short range, captured by these models’ graph scoring mechanisms, are beneficial for solving tasks that require forming a reasoning chain from one or multiple facts to provide factual information. Our AGRAG achieves the best performance on Fact Retrieval and Complex Reasoning tasks. as it can benefit from phrase-level dependencies by generating explicit reasoning chains and directly feeding it to the LLM as guidance.
Lastly, in the text classification task, HippoRAG2 [31] achieves the second-best performance across all metrics, slightly outperformed only by AGRAG. This is because text classification requires capturing correlations between specific text entities and the text chunks to be classified, and both HippoRAG2 and AGRAG explicitly model entity-to-passage relationships in the KG. AGRAG achieves the best performance as its reasoning subgraph spans multiple entity nodes in the KG and explicitly captures the relationships between these entities and the text chunks associated with a given classification label, which can assist LLM classification.
@@ IV-C Efficiency Evaluation
TABLE VI: The ablation experiment result on the two datasets and 4 tasks from GraphRAG-bench [34] benchmark. All experiments were carried out with Qwen2.5-14B-Instruct LLM backbone.
Model Creative Gen. Contextual Sum. Complex Rea. Fact Retrieval Total Efficiency
@@
FS. COV. FS. COV. COV. ROUGE-L Time Token Time Token
AGRAG w. LLM ER 0.4275 0.3650 0.3519 0.5873 0.7553 0.6000 0.3196 0.3330 0.3220 0.3525 244.50 8,917 180.71 5,971
AGRAG w.o. HR 0.5194 0.3433 0.3419 0.5701 0.6771 0.4199 0.3175 0.3240 0.3136 0.3435 200.64 5,615 120.79 5,135
AGRAG w.o. MCMI 0.4882 0.3580 0.3506 0.5749 0.7479 0.6003 0.3242 0.3292 0.3221 0.3587 194.27 7,467 134.50 5,561
AGRAG 0.5134 0.3857 0.3560 0.5888 0.7779 0.6904 0.3186 0.3431 0.3241 0.3572 207.45 7,627 154.42 5,558
To evaluate the efficiency of AGRAG, we compare AGRAG’s total time and token cost with other baselines’ on all 4 dataset tested, the results are shown in Table V. According to the result, AGRAG can also achieve the best time and token efficiency compared with other baselines. It achieves at most 1.66 times faster compared with GraphRAG on Novel dataset, and at least 10.5% accleration compared to HippoRAG2 on Novel dataset. As for token saving, AGRAG also achieves at most 3.69 times token saving compared with GraphRAG on IFS-REL dataset, and save 9.3% tokens at least compared with the previous state-of-the-art model HippoRAG2.
@@ IV-D Ablation Study
To analyze AGRAG’s performance, we conduct an extensive ablation study on GraphRAG-bench tested across 2 datasets and 4 tasks, the ablation models tested is listed as follow:
@@
Firstly, although AGRAG w. LLM ER is more time and token-consuming, it still achieves worse performance compared to AGRAG. Based on the case study in subsection IV-F. This is because the LLM entity extraction approach can produce hallucinated outputs that fail to conform to the expected entity extraction format, leading to decoding errors. In this scenario, current models either skip these entities or text chunks, resulting in incomplete knowledge, or construct graph with the hallucinated outputs, leading to low graph construction quality. This highlights the importance of substituting the LLM entity extraction method to avoid hallucination.
@@
Secondly, AGRAG w.o. HR experiences a significant performance drop on the Contextual Summarization task and in the coverage score of the Creative Generation task. This is because it provides the LLM with only the reasoning path and its associated text chunks—while efficient, this limits the amount of supportive content available and restricts the model’s ability to find sufficient grounded information.
At last, AGRAG w.o. MCMI performs similarly to AGRAG on the Fact Retrieval and Complex Reasoning tasks. This is because its simple, tree-structured reasoning paths are insufficient to support queries from harder tasks, which may require the inclusion of cyclic structures or multiple answer nodes that are difficult to jointly capture within the Steiner tree framework.
IV-E Parameter Sensitivity Analysis
(a) A comparasion between the LLM entity extraction of HippoRAG2 and AGRAG’s statistic-based entity extraction. Compared with LLM entity extraction method, our method extracts more candidate entities, lead to generating more natural and semantically complete triples.
(b) A comparison between generating the Steiner tree and the MCMI for reasoning, where the red edges denote those relations within both the Steiner tree and the MCMI, and the orange edges denote relations solely within the MCMI. In this case, the Steiner tree does not retrieve the usage example and fail to answer the second question.
Figure 5: Two cases of AGRAG w.r.t. its statistics-based entity extraction method and MCMI generation step from GraphRAG-bench’s novel dataset.
In this subsection, we analyse the parameter sensitivity of AGRAG’s only two unique hyperparameters: The maximum n-gram $b$ and the entity extraction threshold $\tau$. As all other hyperparameters of AGRAG also exist in other baselines, we just align them with other baselines for fair comparison. We carry out the analysis on GraphRAG-bench [34], which cover most of the tasks of our experiment.
IV-E1 The maximum n-gram $b$
As introduced in subsection III-B, the maximum n-gram $b$ is used as the entity boundary to control the maximum allowed word length of the extracted entities. Here, we grow a test about the maximum n-gram parameter $b\in\{1,3,5,7\}$, as entity distribution in English is, in general, less than 7 words according to [80]. From the experiment results in Figure 4, AGRAG achieves the overall best performance when the maximum n-gram $b$ is set to 3, this is because most of the entities usually consist of no more than 3 words, according to the discovery of [81, 82, 80].
For Creative Generation task, AGRAG can achieve better performance with larger maximum n-gram, because the correct answer of this task are more flexiable than other tasks that have fixed answers or require answers to cover certain facts. Therefore, the negative impact of noise introduced by high n-gram phrases is overwhelmed by the additional context they provide, leading to better performance. This also highlights AGRAG’s adaptability to different tasks through its flexible graph construction. In our experiments, we set the maximum n-gram $b$ to 3 for all tasks.
@@ IV-E2 The entity extraction threshold $\tau$
As introduced in subsection III-B, in the extraction procedure, we extract words or phrases that have the entity extraction score $\text{ER}(\cdot)>\tau$ as an entity, and we have $0<\text{ER}(\cdot)<1$. Considering the computational cost, we only test the entity extraction threshold $\tau\in\{0.1,0.3,0.5,0.7\}$. According to the experiment results in Figure 4, AGRAG typically achieves its best performance when set to 0.3 or 0.5. This is because a large threshold tends to extract less informative entities, and a small one tends to cause insufficient entity extraction; both can affect the model’s effectiveness. In our experiments, we set the entity extraction threshold $\tau$ to 0.5 for all tasks.
IV-F Case Study
In this subsection, we examine explicit case studies from our testing dataset. Due to space constraints, we focus on example cases related to AGRAG’s statistics-based entity extraction and its MCMI reasoning subgraph generation approach.
Firstly, we compare the entity extraction and triple extraction results of AGRAG with those of HippoRAG2, which employs an LLM-based entity extraction method but shares the same triple extraction process as AGRAG. As shown in Figure 5(a), the statistics-based entity extraction approach can avoid entity extraction failures caused by LLM hallucinations, and extract more candidate entities from the given text chunk, leading to the extraction of more natural and semantically complete triples.
Secondly, we compare AGRAG’s MCMI subgraph with a simple Steiner tree. In Figure 5(b), the Steiner tree is represented by red edges, while the MCMI subgraph includes both red and orange edges. Both reasoning structures successfully answer the first question marked in red in the query. However, the Steiner tree fails to address the second question marked in orange. In contrast, the MCMI subgraph incorporates examples of symbolic usage, such as predilection number for Mayas, enabling the LLM to generate more comprehensive and nuanced answers.
@@ V Conclusion
This paper presents AGRAG, an Advanced Graph-based Retrieval Augmented Generation framework addressing the Inaccurate Graph Construction, Poor Reasoning Ability, and Inadequate Answer issues of current RAG frameworks. AGRAG introduces a statistics-based entity extraction method to replace the LLM entity extraction method applied by current baselines, avoiding hallucination, improving effectiveness, and having better efficiency. By formulating retrieval as a Minimum Cost Maximum Influence (MCMI) subgraph generation problem, AGRAG captures further query-relevant semantic dependencies by allowing more complex graph structure, and the reasoning paths can be more comprehensive and include more query-related graph nodes. By providing explicit reasoning paths to help LLM inference, AGRAG makes the LLM better focus on query-related contexts within the retrieved context. These can lead to better performance on complex tasks that require content summarization. Experiments on real-world datasets demonstrate the effectveness and efficiency of AGRAG over state-of-the-art RAG models.
VI AI-Generated Content Acknowledgement
In this paper, we use AI models, including but not limited to Qwen3-Max and Claude-4.5-sonnet, to polish writing as well as code debugging.

### v2

@@ IV-A4 Hyperparameter and Hardware Settings
For our AGRAG, the length of text chunks $l_{t}$ is set to 256, and token overlap between chunks $l_{o}$ to 32, the number of raw mapped triples $k_{a}$ and the number of hybrid retrieved chunks $k_{r}$ are both set to 5, to align with the setting of HippoRAG2 and GraphRAG-bench. AGRAG’s unique hyperparameter, the entity extraction threshold, and the maximum n-gram, were set to 0.5 and 3, respectively. Please refer to Section IV-E for further parameter sensitivity analysis.
All our experiments were carried out and evaluated with Qwen2.5-14B-Instruct [57] LLM backbone deployed with vLLM [39] if not further illustrated. We apply the Contriever [32] embedding model for HippoRAG2 and AGRAG’s dense passage retrieval procedure, and bge-large-en-v1.5 [75] embedding model for other baselines’ LLM embedding. We apply Qwen-3-Reranker-0.6B LLM for raw triplet filtering. All experiments are conducted on an Intel(R) Xeon(R) Gold 5220R CPU and a single NVIDIA A100-80GB GPU.
@@ IV-B Effectiveness Evaluation
Based on the experiment results in Table III we have observations as follow:
Firstly, in line with recent findings [74, 25, 89], NaiveRAG [22] achieves better performance than all tested walking-based baselines on Contextual Summarization and Creative Generation tasks. This is because, walking-based baselines overlook the global, and coarse-grained information that is essential for effective summarization and creative generation. AGRAG achieves the best performance on these tasks by providing the reasoning subgraph, as well as the hybrid retrieved text chunks to the LLM, offering more comprehensive coarse to fine-grained context for query answering.
Secondly, Our AGRAG achieves the best performance on Fact Retrieval and Complex Reasoning tasks. as it can benefit from phrase-level dependencies by generating explicit reasoning chains and directly feeding it to the LLM as guidance.
Table V: The hallucinated prediction rate on WOS dataset.
Category Model Hallucinated Rate
Pure LLM Qwen-2.5-14B 36.88%
Traditional RAG NaiveRAG 9.96%
L.LLMLingua 21.96%
Tree-based RAG RAPTOR 7.40%
SiReRAG 7.36%
Community-based RAG GraphRAG 6.90%
LightRAG 5.89%
Walking-based RAG HippoRAG2 4.25%
F.GraphRAG 6.06%
AGRAG 0.40%
Lastly, in the text classification task, HippoRAG2 [gutiérrez2025ragmemory] and SiReRAG [82] achieve the second-best performance across all metrics on WOS and IFS-Rel dataset respectively, outperformed by AGRAG. This is because text classification requires capturing correlations between specific text entities and the text chunks to be classified, and all of SiReRAG, HippoRAG2 and AGRAG explicitly model entity-to-passage relationships in the KG. AGRAG achieves the best performance as its reasoning subgraph spans multiple entity nodes in the KG and explicitly captures the relationships between these entities and the text chunks associated with a given classification label, which can assist LLM classification.
Furthermore, Table V presents the hallucinated prediction rate, defined as the percentage of classifications out of the label set. AGRAG achieves the lowest rate by leveraging MCMI subgraphs to provide explicit evidence for retrieved chunks.
@@ IV-C Efficiency Evaluation
Table VI: The ablation experiment result on the two datasets and 4 tasks. All experiments were carried out with Qwen2.5-14B-Instruct LLM, where Time (I / Q) denotes the Index and Query time, Token (I / O) denotes the LLM Input and Output token count.
Model Creative Gen. Contextual Sum. Complex Rea. Fact Retrieval Model Efficiency
@@
FS. ACC. COV. FS. ACC. COV. COV. ACC. COV. ACC. ROG. ACC. ROG. ACC. ROG. ACC. ROG. ACC. Time (I / Q) Token (I / G) Time (I / Q) Token (I / G)
AGRAG w. HR Only 0.491 0.322 0.378 0.303 0.526 0.549 0.739 0.522 0.535 0.583 0.303 0.453 0.334 0.636 0.323 0.487 0.355 0.619 11.7 / 137.9 5,117 / 193 11.4 / 117.9 5,215 / 199
AGRAG w. LLM ER 0.428 0.302 0.365 0.352 0.526 0.587 0.755 0.552 0.600 0.566 0.320 0.465 0.333 0.624 0.322 0.480 0.353 0.613 83.3 / 161.2 8,030 / 887 31.9 / 148.8 5,457 / 414
AGRAG w.o. HR 0.519 0.306 0.343 0.342 0.521 0.570 0.677 0.534 0.420 0.578 0.318 0.434 0.324 0.598 0.314 0.487 0.344 0.615 61.7 / 138.9 6,986 / 741 9.4 / 114.4 4,820 / 315
AGRAG w.o. MCMI 0.488 0.332 0.358 0.351 0.530 0.575 0.748 0.592 0.600 0.588 0.324 0.465 0.329 0.633 0.322 0.470 0.359 0.621 63.9 / 130.4 6,710 / 757 9.5 / 125.0 5,230 / 331
AGRAG w.o. Cle. 0.478 0.311 0.361 0.336 0.517 0.569 0.712 0.597 0.616 0.577 0.301 0.450 0.334 0.628 0.313 0.478 0.354 0.620 65.6 / 140.1 6,988 / 154 9.1 / 142.9 5,252 / 158
AGRAG 0.513 0.339 0.386 0.356 0.598 0.589 0.778 0.643 0.690 0.604 0.319 0.467 0.343 0.665 0.324 0.490 0.357 0.628 65.6 / 142.0 6,988 / 639 9.1 / 145.3 5,252 / 306
Table VII: Low resource experiment effectiveness with Qwen3-4B-Instruct-2507 LLM on 2080Ti GPU, where model with -L denotes low resource setting.
Task Creative Gen. Contextual Sum. Complex Rea. Fact Retrieval
Dataset Novel Medical Novel Medical Novel Medical Novel Medical
SiReRAG-L 0.324 0.412 0.299 0.421 0.758 0.475 0.647 0.673 0.536 0.738 0.260 0.426 0.308 0.698 0.271 0.416 0.343 0.661
HippoRAG2-L 0.433 0.481 0.335 0.425 0.755 0.433 0.696 0.657 0.522 0.731 0.280 0.507 0.298 0.681 0.288 0.513 0.338 0.654
AGRAG-L 0.441 0.519 0.354 0.436 0.777 0.490 0.745 0.706 0.545 0.751 0.293 0.552 0.312 0.739 0.297 0.560 0.351 0.701
Table VIII: Low resource experiment efficiency with Qwen3-4B-Instruct-2507 LLM on 2080Ti GPU, where -L denotes low resource setting.
Dataset Novel Medical
Time (mins) AVG. Token Time (mins) AVG. Token
Metric Index Query In Out Index Query In Out
SiReRAG-L 52.0 183.5 7,898 1,003 18.1 182.3 5,113 827
HippoRAG2-L 55.7 131.2 7,964 1,354 21.8 162.4 5,577 584
AGRAG-L 44.7 132.1 6,625 1,172 14.4 167.3 5,282 534
To evaluate the efficiency of AGRAG, we compare AGRAG’s time and token cost with other baselines’ on all 4 dataset tested, the results are shown in Table IV. According to the result, AGRAG can also achieve the best time and token efficiency compared with other baselines.
@@
Lastly, according to Table VII and Table VIII, where we test AGRAG and 2 other best performing baselines with VLLM [39] deployed Qwen3-4B-Instruct-2507 [78] LLM on two 2080TI GPUs. AGRAG-L still can achieve comparable efficiency under low resource setting, while also achieves better accuracy than normal. This is because the Qwen3-4B is a more novel LLM compared with Qwen2.5-14B, it contains richer general knowledge and can perform better in common-sense reasoning [78]. However, due to the model size limitation, Qwen2.5-14B is still better at context understanding and instruction following, making our normal setting outperform the low-resource setting in other metrics.
@@ IV-D Ablation Study
To analyze AGRAG’s performance, we conduct an extensive ablation study on GraphRAG-bench tested across 2 datasets and 4 tasks, the ablation models tested are listed as follow:
  • AGRAG w. HR Only: Here, we apply only AGRAG’s hybrid retrieval step, without any reasoning-path generation.
@@
  • AGRAG w.o. Cle.: Here, we remove the query aware raw triplet filter procedure (Equation (9)) when graph retrieval.
@@
Firstly, AGRAG w. HR Only achieves the best indexing time efficiency, as it omit all graph related procedures. However, the lack of fine-grained information from graph make it suffer from sub-optimal performance especially on Creative Generation and Contextual Summarization tasks that require summarization to fine-grained information.
Table IX: Breakdown time cost (mins) of AGRAG on 4 datasets.
Dataset
Entity
Extraction
Chunk
Encoding
Triplet
Filter
MCMI
Generation
Hybrid
Retrieval
QA
Novel 62.0 2.8 2.3 3.3 2.4 98.9
Medical 15.1 0.6 2.8 9.1 7.0 72.7
WOS 7.4 0.3 3.4 1.6 0.4 106.7
IFS-Rel 5.6 0.2 1.4 0.3 0.1 98.9
Secondly, although AGRAG w. LLM ER is more time and token-consuming, it still achieves worse performance compared to AGRAG. This is because the LLM entity extraction approach can produce hallucinated outputs that fail to conform to the expected entity extraction format, leading to decoding errors. This highlights the importance of substituting the LLM entity extraction method to avoid hallucination.
Thirdly, AGRAG w.o. HR experiences a significant performance drop on the Contextual Summarization task and in the coverage score of the Creative Generation task. This is because it provides the LLM with only the reasoning path and its associated text chunks—while efficient, this limits the amount of supportive content available and restricts the model’s ability to find sufficient grounded information.
Fourthly, AGRAG w.o. MCMI performs similarly to AGRAG on the Fact Retrieval and Complex Reasoning tasks but consistently underperforms AGRAG on the more challenging Contextual Summarization and Creative Generation tasks. This is because its simple, tree-structured reasoning paths are insufficient to support queries from harder tasks, which may require the inclusion of cyclic structures or multiple answer nodes that are difficult to jointly capture within the Steiner tree framework.
Lastly, AGRAG w.o. Cle. achieves suboptimal performance compared to AGRAG, especially in more complex Creative Generation and Contextual Summarization tasks, which require summarization to graph knowledge. In these scenarios, AGRAG needs the filter procedure to ensure the query-relatedness of retrieved triples, otherwise noise triples are included and affect the model summarization. Considering that omitting the query-related triplet filter can only marginally improve AGRAG’s efficiency (reduce less than 1% of time cost), we consider the triplet cleaning procedure to be worthwhile.
We also show the breakdown time cost in Table IX, where the main time efficiency bottleneck of AGRAG is the LLM QA procedure, which were inevitably suffered by all LLM-based RAG methods, other procedures only lead to a marginal time cost increase. On Novel dataset, although the entity extraction procedure achieve higher time cost, according to Table VI, this cost can be even higher if we apply the LLM based entity extraction approach, because AGRAG w. LLM ER (which applies LLM entity extraction method) achieves higher time and token cost than AGRAG.
IV-E Parameter Sensitivity Analysis
In this subsection, we analyse the parameter sensitivity of AGRAG’s two unique hyperparameters: The maximum n-gram $b$ and the entity extraction threshold $\tau$. As all other hyperparameters of AGRAG also exist in other baselines, we just align them with other baselines for fair comparison. We carry out the analysis on GraphRAG-bench [74], which cover most of the tasks of our experiment.
IV-E1 The maximum n-gram $b$
As introduced in Section III-B, the maximum n-gram $b$ is used as the entity boundary to control the maximum allowed word length of the extracted entities. Here, we conduct a test about the maximum n-gram parameter $b\in\{1,3,5,7\}$, as entity distribution in English is, in general, less than 7 words according to [61]. From the experiment results in Figure 3, AGRAG achieves the overall best performance when the maximum n-gram $b$ is set to 3, this is because most of the entities usually consist of no more than 3 words, according to the discovery of [90, 21, 61].
For Creative Generation task, AGRAG can achieve better performance with larger maximum n-gram, because the correct answer of this task is more flexible than other tasks that have fixed answers or require answers to cover certain facts. Therefore, the negative impact of noise introduced by high n-gram phrases is overwhelmed by the additional context they provide, leading to better performance. This also highlights AGRAG’s adaptability to different tasks through its flexible graph construction. In our experiments, we set the maximum n-gram $b$ to 3 for all tasks.
@@ IV-E2 The entity extraction threshold $\tau$
As introduced in Section III-B, in the extraction procedure, we extract words or phrases that have the entity extraction score $\text{ER}(\cdot)>\tau$ as an entity, and we have $0<\text{ER}(\cdot)<1$. Considering the computational cost, we only test the entity extraction threshold $\tau\in\{0.1,0.3,0.5,0.7\}$. According to the experiment results in Figure 3, AGRAG typically achieves its best performance when set to 0.3 or 0.5. This is because a large threshold tends to extract less informative entities, and a small one tends to cause insufficient entity extraction; both can affect the model’s effectiveness. In our experiments, we set the entity extraction threshold $\tau$ to 0.5 for all tasks.
@@ V Conclusion
This paper presents AGRAG, an Advanced Graph-based Retrieval Augmented Generation framework addressing the Inaccurate Graph Construction, Poor Reasoning Ability, and Inadequate Answer issues of current RAG frameworks. By providing explicit reasoning paths to help LLM inference, AGRAG makes the LLM better focus on query-related contexts within the retrieved context. Experiments on real-world datasets demonstrate the effectiveness and efficiency of AGRAG over state-of-the-art RAG models.
VI Acknowledgements
Lei Chen is supported by National Key Research and Development Program of China Grant No. 2023YFF0725100, National Science Foundation of China under Grant No. U22B2060, Guangdong-Hong Kong Technology Innovation Joint Funding Scheme Project No. 2024A0505040012, AOE Project AoE/E-603/18, Theme-based project TRS T41-603/20R, CRF Project C2004-21G, Key Areas Special Project of Guangdong Provincial Universities 2024ZDZX1006, Guangdong Province Science and Technology Plan Project 2023A0505030011, HKUST(GZ) CMCC(Guangzhou Branch) Metaverse Joint Innovation Lab under Grant No. P00659, Hong Kong ITC TC-SKLCRCC26EG01, ITF grant PRP/004/22FX, Zhujiang scholar program 2021JC02X170, HKUST Webank joint research lab. Haoyang Li’s work is partially supported by grant No. PolyU P0052504, PolyU-Huawei P0053707, PolyU-Minshang P0057770, PolyU TDG25-28/TBP/11-IICA.
VII AI-Generated Content Acknowledgement
In this paper, we use AI models: Qwen3-Max and Gemini3, to polish writing as well as code debugging.